# Arbitrarily-Conditioned Multi-Functional Diffusion for Multi-Physics Emulation

### v1

@@ 3.3 Efficient Training And Sampling
To address this challenge, we use a multiplicative kernel to model the noise function. Given the input dimension $D$, we define the kernel as
@@
Notably, the widely used Square Exponential (SE) kernel already exhibits this structure. We position the sampling locations for each target function $f^{k}_{0}(\cdot)$ on an $m_{1}\times\ldots\times m_{D}$ mesh, denoted as ${\mathcal{Z}}_{k}=\bm{\gamma}_{1}\times\ldots\times\bm{\gamma}_{D}$, where $\times$ denotes the Cartesian product, and each $\bm{\gamma}_{d}$ ($1\leq d\leq D$) comprises the $m_{d}$ input locations within dimension $d$.
@@
${\bf F}^{s}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left({\bf F}_{t}^{s}-\frac{\beta_{t}}{\sqrt{1-\widehat{\alpha}_{t}}}\overline{{\bm{\xi}}}^{s}_{t}\right)+\widehat{{\beta}}_{t}\overline{{\bm{\epsilon}}}$
@@
where we reshape $\bm{\eta}$ into an $m_{1}\times\ldots\times m_{D}$ tensor $\Pi$, and ×$\times_{k}$ is the tensor-matrix product along mode $k$. The time complexity of this operation becomes ${\mathcal{O}}(\overline{m}(\sum_{d}m_{d}))$. Overall, this approach avoids computing the full covariance matrix ${\bf K}$ for all the mesh points along with its Cholesky decomposition, which would require ${\mathcal{O}}(\overline{m}^{3})$ time and ${\mathcal{O}}(\overline{m}^{2})$ space complexity. Instead we only compute the local kernel matrix for each input dimension. For example, when using a $128\times 128$ mesh, the full covariance matrix would be $128^{2}\times 128^{2}$, which is highly expensive to compute. In contrast, our approach requires the computation of only two local kernel matrices, each of size $128\times 128$. Our approach further uses the Tucker product to compute the sample (see (13)), thereby avoiding the computationally expensive Kronecker product. As a result, the overall time and space complexity is reduced to ${\mathcal{O}}(\sum_{d}m_{d}^{3}+\overline{m}(\sum_{d}m_{d}))$ and ${\mathcal{O}}(\sum_{d}m_{d}^{2}+\overline{m})$, respectively.
In cases where the data is irregularly and sparsely sampled — often the case with practical observation or measurement data — we employ a GP model to interpolate the function values onto a grid. This allows us to efficiently train and sample using the aforementioned approach. Empirically, we have found that this straightforward strategy demonstrates robust performance, often surpassing alternative approaches such as attention mechanisms (see Section 5.1). Finally, we summarize the training and generation of ACM-FD in Algorithm 1 and 2, respectively.
@@ 4 RELATED WORK
Generative models have been studies for a long time. One class is the likelihood-based approach, which uses neural networks to approximate the probability density function, and maximize the likelihood of the observed data under the model. This category includes auto-regressive models (Larochelle and Murray, 2011; Germain et al., 2015; Van Den Oord et al., 2016), variational auto-encoders (Kingma and Welling, 2013; Rezende et al., 2014), normalization flows (Dinh et al., 2014, 2016; Papamakarios et al., 2017), and energy-based models LeCun et al. (2006); Gutmann and Hyvärinen (2010); Song and Kingma (2021), and others. A second class is implicit generative models (Mohamed and Lakshminarayanan, 2016), which transforms a random variable drawn from a standard distribution (e.g., Gaussian) into a new variable, typically, via neural networks, resulting in a complex implicit density model. Typical examples of such methods include Generative Adversarial Networks (GAN) and their variants Goodfellow et al. (2014); Salimans et al. (2016); Arjovsky et al. (2017).
The recent DDPM work (Ho et al., 2020) is a break-through, which specifies a probability-path (diffusion process) that progressively converts data instances into noise and then learns to reverse this path. Numerous subsequent works have built upon this direction, such as score-based diffusion via SDEs (Song et al., 2021b), denoising diffusion implicit models (DDIM) (Song et al., 2021a), flow-matching (Lipman et al., 2022; Klein et al., 2023) fast sampling (Salimans and Ho, 2022; Kong and Ping, 2021; Zhang and Chen, 2023; Shih et al., 2023), latent diffusion (Rombach et al., 2022; Avrahami et al., 2023), and others. The most recent work (Zhang and Wonka, 2024) proposed functional diffusion under the DDIM framework. The key difference is that it focuses on single-function generation and does not address multi-functional scenarios or diverse conditional sampling tasks, as we do. Another notable recent work by Kerrigan et al. (2024) extends the flow-matching framework, based on neural Ordinary Differential Equations (Chen et al., 2018), for functional generation. In addition to being based on different generative modeling frameworks, our work primarily differs in that we train a single model to handle a diverse set of conditioned and unconditioned functional sampling tasks, whereas (Kerrigan et al., 2024) focuses on a single task.
@@
Neural operator (NOs) are emerging as a promient research area in ML-based surrogate models. Their goal is to learn function-to-function mappings from data, primarily for estimating PDE operators from simulation data and performing forward predictions of PDE solutions given new inputs. Important works include Fourier Neural Operators (FNO) (Li et al., 2022a) and Deep Operator Net (DONet) (Lu et al., 2021), which have demonstrated promising prediction accuracy across various benchmark PDEs (Lu et al., 2022), such as Burger’s equation, Darcy-flow and Naiver-Stoke (NS) equations. Many NO models have been developed based on FNO and DONet. For example, Gupta et al. (2021) contributed with multiwavelet transformations for the operator’s kernel. Lu et al. (2022) proposed POD-DONet to enhance the stability of DONet by replacing the trunk net with POD (PCA) bases constructed from data. A recent line of research attempts to build neural operators using attention/transformer architectures, such as (Cao, 2021; Li et al., 2022b; Hao et al., 2023). Additionally, recent advances in kernel operator learning strategies have been made by Long et al. (2022) and Batlle et al. (2023). A recent review of NOs can be found in (Kovachki et al., 2023).
@@ 5 EXPERIMENTS
For evaluation, we considered three fundamental multi-physics systems: Darcy Flow: a single-phase 2D Darcy flow system that involves three functions: the permeability field $a$, the flow pressure $u$, and the source term $f$. Convection Diffusion: a 1D convection-diffusion system that involves three spatial-temporal functions: the scalar field of the quantity of interest $u$, the velocity field $v$, and the source function $s$. Torus Fluid: a viscous, incompressible fluid in the unit torus, for which we are interested in seven functions, including the source function $f$, the initial vorticity field $w_{0}$, and the vortocity fields $w_{1},\ldots,w_{5}$ at five different time steps $t=2,4,6,8,10$.
@@
Table 1: Relative $L_{2}$ Error For Various Prediction Tasks, With Training And Test Functions Sampled On Regular Meshes. The Results Were Averaged Over Five runs. D-F, C-D And T-F Are Short For Darcy Fow, Convection Diffusion, And Torus Fluid Systems.
Dataset Task(s) ACM-FD FNO GNOT DON PODDON
D-F $f,u$ to $a$ 1.32e-02 (2.18e-04) 1.88e-02 (1.66e-04) 1.35e-01 (6.57e-05) 2.38e-02 (3.45e-04) 1.56e-02 (1.44e-04)
$a,u$ to $f$ 1.59e-02 (1.59e-04) 2.37e-02 (1.87e-04) 1.00e+00 (0.00e+00) 3.76e-02 (7.75e-04) 4.63e-02 (1.21e-03)
$a,f$ to $u$ 1.75e-02 (4.16e-04) 6.29e-02 (4.18e-04) 6.09e-01 (2.40e-01) 6.05e-02 (7.17e-04) 6.80e-02 (2.16e-04)
$u$ to $a$ 3.91e-02 (7.08e-04) 5.57e-02 (4.16e-04) 1.35e-01 (1.99e-04) 5.08e-02 (5.91e-04) 4.09e-02 (4.18e-04)
$u$ to $f$ 3.98e-02 (6.45e-04) 5.50e-02 (5.47e-04) 9.99e-01 (7.48e-04) 6.46e-02 (1.13e-04) 6.37e-02 (1.03e-03)
C-D $s,u$ to $v$ 2.17e-02 (4.53e-04) 4.50e-02 (3.89e-04) 3.26e-02 (3.41e-03) 3.64e-02 (5.07e-04) 4.62e-02 (2.00e-04)
$v,u$ to $s$ 5.45e-02 (1.40e-03) 7.93e-02 (8.48e-04) 1.22e-01 (1.91e-03) 7.04e-02 (7.53e-04) 7.57e-02 (4.09e-04)
$v,s$ to $u$ 1.60e-02 (2.15e-04) 7.26e-02 (2.16e-04) 5.80e-03 (1.51e-04) 7.86e-02 (7.42e-04) 1.72e-01 (1.17e-03)
$u$ to $v$ 2.66e-02 (3.08e-04) 5.90e-02 (8.22e-04) 6.69e-02 (3.66e-03) 4.55e-02 (6.09e-04) 5.38e-02 (5.29e-04)
$u$ to $s$ 6.06e-02 (2.54e-04) 1.16e-01 (5.63e-04) 1.85e-01 (2.84e-03) 9.65e-02 (5.52e-04) 1.03e-01 (1.29e-03)
T-F $w_{0},w_{5}$ to $w_{1}$ 2.73e-02 (4.78e-03) 1.28e-02 (2.38e-04) 2.40e-02 (8.74e-04) 6.32e-02 (2.72e-04) 6.06e-02 (2.91e-04)
$w_{0},w_{5}$ to $w_{2}$ 2.43e-02 (1.60e-03) 2.08e-02 (9.80e-05) 4.00e-02 (5.92e-04) 7.69e-02 (4.41e-04) 7.71e-02 (1.63e-04)
$w_{0},w_{5}$ to $w_{3}$ 2.43e-02 (3.17e-03) 2.33e-02 (1.83e-04) 4.74e-02 (1.23e-03) 7.34e-02 (2.88e-04) 7.38e-02 (2.92e-04)
$w_{0},w_{5}$ to $w_{4}$ 1.68e-02 (1.81e-03) 1.41e-02 (1.17e-04) 3.95e-02 (6.73e-04) 5.57e-02 (1.73e-04) 5.38e-02 (2.18e-04)
$w_{0},w_{5}$ to $f$ 1.63e-02 (1.49e-03) 1.79e-02 (3.04e-04) 5.91e-02 (4.01e-03) 4.77e-02 (5.56e-04) 4.94e-02 (9.02e-04)
$w_{0},f$ to $w_{1}$ 3.10e-02 (4.08e-03) 9.68e-03 (3.22e-04) 2.09e-02 (3.62e-04) 6.08e-02 (3.14e-04) 5.54e-02 (6.45e-04)
$w_{0},f$ to $w_{2}$ 3.28e-02 (4.79e-03) 1.70e-02 (3.51e-04) 4.15e-02 (8.21e-04) 7.73e-02 (6.18e-04) 7.40e-02 (2.01e-04)
$w_{0},f$ to $w_{3}$ 3.49e-02 (2.38e-03) 2.38e-02 (8.37e-05) 5.61e-02 (8.23e-04) 8.82e-02 (4.45e-04) 8.60e-02 (5.16e-04)
$w_{0},f$ to $w_{4}$ 3.34e-02 (3.87e-03) 3.10e-02 (1.26e-04) 6.97e-02 (1.62e-03) 1.02e-01 (7.28e-04) 9.74e-02 (4.21e-04)
$w_{0},f$ to $w_{5}$ 3.26e-02 (2.13e-03) 3.81e-02 (2.01e-04) 8.35e-02 (7.33e-04) 1.21e-01 (8.20e-04) 1.16e-01 (8.07e-04)
@@ 5.1 Predictive Performance Across Various Tasks
Methods. We compared with the following popular and state-of-the-art operator learning methods: (1) Fourier Neural Operator (FNO) (Li et al., 2020b), which performs channel lifting, followed by a series of Fourier layers that execute linear functional transformations using the fast Fourier transform and then apply nonlinear transformations, ultimately culminating in a projection layer that produces the prediction. (2) Deep Operator Net (DON) (Lu et al., 2021), which employs a branch-net and a trunk-net to extract representations of the input functions and querying locations, producing the prediction by the dot product between the two representations. (3) PODDON (Lu et al., 2022), a variant of DON where the trunk-net is replaced by the POD (PCA) bases extracted from the training data. (4) GNOT (Hao et al., 2023), a transformer-based neural operator that uses cross-attention layers to aggregate multiple input functions’ information for prediction. We utilized the original implementation of each competing method. Our method, ACM-FD, was implemented with PyTorch. For GP noise function sampling and interpolation of irregularly sampled functions, we used the Square-Exponential (SE) kernel. In the experiments, we used FNO to construct our denoising network $\Phi_{\Theta}$. Thereby, our comparison with FNO can exclude the factors from the architecture design.
@@
Setting. We employed two evaluation settings for each task. In the first setting, all functions are sampled on a regular $64\times 64$ grid, while the second setting samples each function from random input locations within the domain, with the number of samples varying from 1,200 to 1,500. To enable FNO and PODDON to train with irregularly sampled data, we used scipy.interpolate.griddata routine to interpolate the sampled functions onto a $64\times 64$ grid. Note that DON and GNOT can directly process irregularly sampled data. For both the regularly and irregularly sampled settings, we utilized 1,000 instances for training, 100 instances for validation, and 200 instances for testing. Hyperparameter tuning was performed using the validation set. The details are provided in Appendix Section B. Following the evaluation procedure in (Lu et al., 2022), for each method, we selected the optimal hyperparameters, and then conducted stochastic training for five times, reporting the average relative $L_{2}$ test error along with the standard deviation. The results for the regularly-sampled case and irregularly-sampled case are presented in Table 1 and 2, respectively.
Result. As we can see, ACM-FD consistently achieved top-tier predictive performance across all tasks. In the regularly-sampled setting, ACM-FD outperforms FNO in all tasks for both Darcy Flow and Convection Diffusion. In the Torus Fluid system , while ACM-FD ranks second, its performance remains close to that of FNO. In the irregularly sampled setting, ACM-FD consistently outperforms all competing methods, often by a large margin. It is worth noting that GNOT is a transformer based approach, and by using the cross-attention layers, GNOT can flexibly handle irregularly sampled data, without the need to conduct explicit interpolation. However, GNOT still consistently performs worse than ACM-FD across all the tasks, despite being specifically tuned for each task. A similar observation holds for DON. These results demonstrate the superior performance and versatility of ACM-FD.

### v2

@@ 3.3 Efficient Training and Sampling
To address this challenge, we use a multiplicative kernel to model the noise function. Given the input dimension $D$, we construct the kernel with the following form,
@@
Notably, the widely used Square Exponential (SE) kernel already exhibits this structure: $\kappa_{\text{SE}}({\bf z},{\bf z}^{\prime})=\exp(-\|{\bf z}-{\bf z}^{\prime}\|^{2}/l^{2})$. We position the sampling locations for each target function $f^{k}_{0}(\cdot)$ on an $m_{1}\times\ldots\times m_{D}$ mesh, denoted as ${\mathcal{Z}}_{k}=\bm{\gamma}_{1}\times\ldots\times\bm{\gamma}_{D}$, where $\times$ denotes the Cartesian product, and each $\bm{\gamma}_{d}$ ($1\leq d\leq D$) comprises the $m_{d}$ input locations within dimension $d$.
@@
${\bf F}^{s}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left({\bf F}_{t}^{s}-\frac{\beta_{t}}{\sqrt{1-\widehat{\alpha}_{t}}}\overline{{\bm{\xi}}}^{s}_{t}\right)+\sqrt{\widehat{{\beta}}_{t}}\overline{{\bm{\epsilon}}}$
@@
where we reshape $\bm{\eta}$ into an $m_{1}\times\ldots\times m_{D}$ tensor $\Pi$, and ×$\times_{k}$ is the tensor-matrix product along mode $k$. The time complexity of this operation becomes ${\mathcal{O}}(\overline{m}(\sum_{d}m_{d}))$. Overall, this approach avoids computing the full covariance matrix ${\bf K}$ for all the mesh points along with its Cholesky decomposition, which would require ${\mathcal{O}}(\overline{m}^{3})$ time and ${\mathcal{O}}(\overline{m}^{2})$ space complexity. Instead we only compute the local kernel matrix for each input dimension. For example, when using a $128\times 128$ mesh, the full covariance matrix would be $128^{2}\times 128^{2}$, which is highly expensive to compute. In contrast, our approach requires the computation of only two local kernel matrices, each of size $128\times 128$. Our approach further uses the Tucker product to compute the noise function sample (see (13)), avoiding explicitly computing the expensive Kronecker product. As a result, the overall time and space complexity is reduced to ${\mathcal{O}}(\sum_{d}m_{d}^{3}+\overline{m}(\sum_{d}m_{d}))$ and ${\mathcal{O}}(\sum_{d}m_{d}^{2}+\overline{m})$, respectively. This enables efficient training and generation of our model. Finally, we summarize our method in Algorithm 1 and 2.
@@ 4 Related Work
Generative modeling is a fundamental topic in machine learning. The DDPM (Ho et al., 2020) is recent a break-through, and numerous subsequent works have expanded upon this direction, further advancing the field, such as score-based diffusion via SDEs (Song et al., 2021b), denoising diffusion implicit models (DDIM) (Song et al., 2021a), and flow-matching (Lipman et al., 2022; Klein et al., 2023). Recent works have explored diffusion models for inverse problems, such as (Chung et al., 2023; Wang et al., 2023a; Song et al., 2023). However, these approaches mainly train an unconditional diffusion model, and then perform conditional sampling at the inference stage.
A few recent works have explored diffusion models for function generation, such as (Kerrigan et al., 2023) based on DDPM, (Lim et al., 2023; Franzese et al., 2024) leveraging score- or SDE-based diffusion, (Zhang & Wonka, 2024) using DDIM, and (Kerrigan et al., 2024) with flow matching. Our approach, which incorporates a GP noise function into the DDPM framework, is conceptually similar to (Kerrigan et al., 2023), where a GP is treated as a special case of a Gaussian random element in their formulation. However, the key distinction lies in the scope of our work. Unlike prior methods focused on single-function generation, our approach is designed for multi-functional scenarios, simultaneously addressing diverse single- and multi-functional generation tasks conditioned on varying quantities and/or functions during inference. To achieve this, we propose an arbitrarily conditioned multi-functional diffusion framework that integrates a random masking strategy, a zero-regularized denoising loss, and an efficient training and sampling method. Our work is related to (Gloeckler et al., 2024), which leverages score-based diffusion for simulation and also uses a random masking strategy to fulfill flexible conditional sampling. But (Gloeckler et al., 2024) does not enforce zero regularization on the score model’s predictions for the conditioned part. In contrast, our method incorporates zero regularization in the loss function to prevent the denoising network from producing excessive artificial noise to the conditioned part, thereby enhancing stability. A detailed comparison is provided in Section 5. Another related work is (Wang et al., 2023b), which introduces a score-based diffusion model for multi-fidelity simulation. However, the approach is not formulated as function generation and is limited to a single prediction task.
@@
Neural operator (NOs) aim to learn function-to-function mappings from data, primarily for estimating PDE operators from simulation data and performing forward predictions of PDE solutions given new inputs (e.g., parameters, initial and/or boundary conditions, source terms). Important works include Fourier Neural Operators (FNO) (Li et al., 2022) and Deep Operator Net (DONet) (Lu et al., 2021), which have demonstrated promising prediction accuracy across various benchmark PDEs (Lu et al., 2022). Many NO models have been developed based on FNO and DONet, such as (Gupta et al., 2021; Wen et al., 2022; Lu et al., 2022; Tran et al., 2023). Li et al. (2024) developed an active learning method to query multi-resolution data for enhancing FNO training while reducing the data cost. Recent advances in kernel operator learning have been made by Long et al. (2022),  Batlle et al. (2023), and Lowery et al. (2024).
@@ 5 Experiments
For evaluation, we considered four fundamental multi-physics systems:
  • Darcy Flow (D-F): a single-phase 2D Darcy flow system that involves three functions: the permeability field $a$, the flow pressure $u$, and the source term $f$.
  • Convection Diffusion (C-D): a 1D convection-diffusion system that involves three spatial-temporal functions: the scalar field of the quantity of interest $u$, the velocity field $v$, and the source function $s$.
  • Diffusion Reaction (D-R): a 2D diffusion-reaction system, for which we are interested in four spatial functions: $f_{1}$ and $f_{2}$, which represent the activator and inhibitor functions at time $t=2.5$, and $u_{1}$ and $u_{2}$, denoting the activator and inhibitor at $t=5.0$.
  • Torus Fluid (T-F): a viscous, incompressible fluid in the unit torus, for which we are interested in seven functions, including the source function $f$, the initial vorticity field $w_{0}$, and the vortocity fields $w_{1},\ldots,w_{5}$ at five different time steps $t=2,4,6,8,10$.
@@
Table 1: Relative $L_{2}$ error for various prediction tasks. The results were averaged over five runs.
Dataset Task(s) ACM-FD FNO GNOT DON Simformer
D-F $f,u$ to $a$ 1.32e-02 (2.18e-04) 1.88e-02 (1.66e-04) 1.35e-01 (6.57e-05) 2.38e-02 (3.45e-04) 1.18e-01 (3.00e-03)
$a,u$ to $f$ 1.59e-02 (1.59e-04) 2.37e-02 (1.87e-04) 1.00e+00 (0.00e+00) 3.76e-02 (7.75e-04) 4.11e-02 (2.87e-03)
$a,f$ to $u$ 1.75e-02 (4.16e-04) 6.29e-02 (4.18e-04) 6.09e-01 (2.40e-01) 6.05e-02 (7.17e-04) 4.04e-02 (5.17e-03)
$u$ to $a$ 3.91e-02 (7.08e-04) 5.57e-02 (4.16e-04) 1.35e-01 (1.99e-04) 5.08e-02 (5.91e-04) 1.44e-01 (4.23e-03)
$u$ to $f$ 3.98e-02 (6.45e-04) 5.50e-02 (5.47e-04) 9.99e-01 (7.48e-04) 6.46e-02 (1.13e-04) 1.06e-01 (3.98e-03)
C-D $s,u$ to $v$ 2.17e-02 (4.53e-04) 4.50e-02 (3.89e-04) 3.26e-02 (3.41e-03) 3.64e-02 (5.07e-04) 3.96e-01 (4.79e-02)
$v,u$ to $s$ 5.45e-02 (1.40e-03) 7.93e-02 (8.48e-04) 1.22e-01 (1.91e-03) 7.04e-02 (7.53e-04) 5.76e-02 (7.10e-02)
$v,s$ to $u$ 1.60e-02 (2.15e-04) 7.26e-02 (2.16e-04) 5.80e-03 (1.51e-04) 7.86e-02 (7.42e-04) 1.03e-01 (1.95e-02)
$u$ to $v$ 2.66e-02 (3.08e-04) 5.90e-02 (8.22e-04) 6.69e-02 (3.66e-03) 4.55e-02 (6.09e-04) 5.108-01 (7.56e-02)
$u$ to $s$ 6.06e-02 (2.54e-04) 1.16e-01 (5.63e-04) 1.85e-01 (2.84e-03) 9.65e-02 (5.52e-04) 9.21e-01 (1.00e-01)
D-R $f_{1},u_{1}$ to $f_{2}$ 1.44e-02 (8.96e-04) 1.07e-02 (1.92e-04) 4.53e-01 (4.34e-02) 2.93e-01 (1.29e-03) 3.39e-02 (2.97e-03)
$f_{1},u_{1}$ to $u_{2}$ 1.59e-02 (3.68e-04) 2.02e-02 (2.42e-04) 3.91e-01 (1.86e-02) 2.03e-01 (2.22e-03) 3.67e-02 (2.36e-03)
$f_{2},u_{2}$ to $f_{1}$ 4.10e-02 (8.93e-04) 5.52e-02 (3.01e-03) 6.53e-01 (2.04e-02) 4.24e-01 (9.26e-04) 1.21e-01 (3.11e-03)
$f_{2},u_{2}$ to $u_{1}$ 5.86e-02 (3.43e-04) 7.82e-02 (1.29e-04) 4.88e-01 (2.92e-02) 2.98e-01 (2.61e-03) 1.01e-01 (2.70e-03)
T-F $w_{0},w_{5}$ to $w_{1}$ 2.73e-02 (4.78e-03) 1.28e-02 (2.38e-04) 2.40e-02 (8.74e-04) 6.32e-02 (2.72e-04) 6.14e-02 (2.44e-03)
$w_{0},w_{5}$ to $w_{2}$ 2.43e-02 (1.60e-03) 2.08e-02 (9.80e-05) 4.00e-02 (5.92e-04) 7.69e-02 (4.41e-04) 6.99e-02 (2.18e-03)
$w_{0},w_{5}$ to $w_{3}$ 2.43e-02 (3.17e-03) 2.33e-02 (1.83e-04) 4.74e-02 (1.23e-03) 7.34e-02 (2.88e-04) 8.34e-02 (2.60e-03)
$w_{0},w_{5}$ to $w_{4}$ 1.68e-02 (1.81e-03) 1.41e-02 (1.17e-04) 3.95e-02 (6.73e-04) 5.57e-02 (1.73e-04) 9.75e-02 (3.93e-03)
$w_{0},w_{5}$ to $f$ 1.63e-02 (1.49e-03) 1.79e-02 (3.04e-04) 5.91e-02 (4.01e-03) 4.77e-02 (5.56e-04) 1.14e-01 (4.00e-03)
$w_{0},f$ to $w_{1}$ 3.10e-02 (4.08e-03) 9.68e-03 (3.22e-04) 2.09e-02 (3.62e-04) 6.08e-02 (3.14e-04) 6.06e-02 (2.03e-03)
$w_{0},f$ to $w_{2}$ 3.28e-02 (4.79e-03) 1.70e-02 (3.51e-04) 4.15e-02 (8.21e-04) 7.73e-02 (6.18e-04) 6.18e-02 (1.02e-03)
$w_{0},f$ to $w_{3}$ 3.49e-02 (2.38e-03) 2.38e-02 (8.37e-05) 5.61e-02 (8.23e-04) 8.82e-02 (4.45e-04) 5.67e-02 (1.83e-03)
$w_{0},f$ to $w_{4}$ 3.34e-02 (3.87e-03) 3.10e-02 (1.26e-04) 6.97e-02 (1.62e-03) 1.02e-01 (7.28e-04) 4.10e-02 (1.98e-03)
$w_{0},f$ to $w_{5}$ 3.26e-02 (2.13e-03) 3.81e-02 (2.01e-04) 8.35e-02 (7.33e-04) 1.21e-01 (8.20e-04) 1.18e-01 (4.15e-03)
@@ 5.1 Predictive Performance Across Various Tasks
We compared with the following popular and state-of-the-art operator learning methods: (1) Fourier Neural Operator (FNO) (Li et al., 2020b), which performs channel lifting, followed by a series of Fourier layers that execute linear functional transformations using the fast Fourier transform and then apply nonlinear transformations, ultimately culminating in a projection layer that produces the prediction. (2) Deep Operator Net (DON) (Lu et al., 2021), which employs a branch-net and a trunk-net to extract representations of the input functions and querying locations, producing the prediction by the dot product between the two representations. (3) PODDON (Lu et al., 2022), a variant of DON where the trunk-net is replaced by the POD (PCA) bases extracted from the training data. (4) GNOT (Hao et al., 2023), a transformer-based neural operator that uses cross-attention layers to aggregate multiple input functions’ information for prediction. We utilized the original implementation of each competing method. In addition, we compared with (5) Simformer (Gloeckler et al., 2024), an attention based emulator which also uses a random masking strategy to fulfill flexible conditional generation. However, Simformer does not fit zero values for the conditioned part as in our model; see Equations (10) and (11). The original Simfomer was developed with score-based diffusion (Song et al., 2021b). For a fair comparison, we adapted Simformer to the DDPM framework, and reimplemented it using PyTorch. Our method, ACM-FD, was implemented with PyTorch as well. For GP noise function sampling, we used the Square-Exponential (SE) kernel. In all the experiments, we used FNO to construct our denoising network $\Phi_{\Theta}$. Thereby, our comparison with FNO can exclude the factors from the architecture design. We combined two types of random masking: one that masks individual observed values within each function, and the other that masks entire functions, both with a masking probability of 0.5. The corresponding denoising losses for these two masking strategies (see (10)) are summed to form the final loss used for training the model.
@@
We utilized 1,000 instances for training, 100 instances for validation, and 200 instances for testing. Hyperparameter tuning was performed using the validation set. The details are provided in Appendix Section B. Following the evaluation procedure in (Lu et al., 2022), for each method, we selected the optimal hyperparameters, and then conducted stochastic training for five times, reporting the average relative $L_{2}$ test error along with the standard deviation. The results are presented in Table 1. Due to the space limit, the comparison results with PODDON are provided in Appendix Table 5.
As we can see, ACM-FD consistently achieves top-tier predictive performance across all the tasks. In Darcy Flow (D-F), Convection Diffusion (C-D) and Diffusion Reaction (D-R), ACM-FD outperforms FNO in all tasks except for the $f_{1},u_{1}$ to $f_{2}$ mapping. In the Torus Fluid (T-F) system, while ACM-FD ranks second, its performance remains close to that of FNO. Furthermore, the performance of ACM-FD is consistently better than Simformer and other neural operators, including GNOT, DON and PODDON (see Table 5 in the Appendix). Despite being specifically optimized for each task, those neural operators are constantly worse than our method, underscoring the advantage of training a single, versatile model. The observed improvement of Simformer over GNOT in many cases (both are based on transformer architectures) also implied the effectiveness of this strategy.